An infinite number of identical objects each of mass m are placed at a distance of 1m.,2m.,3m.,...resp.The gravitatioal potential due to these masses at O is
1]-GM
2]-2GM
3]-4GM
4]-8GM

Grade:11

2 Answers

Nishant Vora IIT Patna
7 years ago
- \frac{GM}{1 }-\frac{GM}{2 }-\frac{GM}{3}+........

Piyush Kumar Behera
7 years ago
The question is not clearly stated  What is the point O please clearly mention it and do please repost the question the question!!!